OpenStudy (anonymous):

y-5=-3/4(x+5) in standard form

OpenStudy (anonymous):

Is it the equation of the line?

OpenStudy (anonymous):

ax+by=c is the formula

OpenStudy (anonymous):

idk it just says in standard form its either a) 3x-4y=-10 b)3x+4y=35 c)3x+4y=5 d) 3x-4y=0

OpenStudy (imstuck):

Does that mean you want to solve this for y? First distribute the -3/4 into the (x + 5) to get -3/4 x - 15/4. Now add 5 to both sides. Of course, since you have a fraction on one side, you need to change the 5 into a fraction with a 4 in the denominator. That would be equivalent to 20/4. So we have - 3/4x - 15/4 + 20/4 which is y = -3/4 x + 5/4. That is the slope-intercept form. If you need another form, there are others, as well. ax + by + c = 0 or ax + by = c...

OpenStudy (anonymous):

thank u i couldnt figure out how to do that one problem

OpenStudy (anonymous):

in the standard form x and y are in left hand side and the constant is on the right side. So, simplify and bring x and y terms to left and send the constant terms to the right.

OpenStudy (anonymous):

Is this your real equation \[y-5=-\frac{ 3 }{ 4 }(x+5)\]

OpenStudy (anonymous):

because that would mean that \[-3(x+5) = 4(y-5) => -3x -15 = 4y - 20 \]